OpenStudy (anonymous):

need help on this one!!! A certain mountain has an elevation of 19,807 feet. In 1925, the glacier on this peak covered 5 acres. By 2004 this glacier had melted to only 1 acre. A) Assume that this glacier melted at a constant rate each year. Find the yearly rate. B) What is the equation that gives the acreage of the glacier t years after 1925?

OpenStudy (anonymous):

I tried sending you a message, bu it won't let me/

OpenStudy (anonymous):

rate is found by taking the amount melted, and dividing by the number of years it took to melt

OpenStudy (anonymous):

it went from 5 to 1 so it lost 4 acres it took form 1925 to 2004 so it took 79 years to meld 4 acres, rate is \(-\frac{4}{79}\)

OpenStudy (anonymous):

*melt

OpenStudy (anonymous):

5-1=4 acres melted, then divide that by ...79 years took to melt 4 / 79 = 0.05 acres melted

OpenStudy (anonymous):

it is negative, because it is a loss of 4 acres

OpenStudy (anonymous):

-0.05 acres melted per year?

OpenStudy (anonymous):

if you round, yes if not, it is \(-\frac{4}{79}\)

OpenStudy (anonymous):

then since you had 5 acres when you started counting in 1925 the formula would be \[y=-\frac{4}{79}x+5\] where \(x\) is years after 1925

OpenStudy (anonymous):

or rather \[y=-\frac{4}{79}t+5\]

OpenStudy (anonymous):

since they are using \(t\) and not \(x\)

OpenStudy (anonymous):

ok I was confused for a minute LOL

OpenStudy (anonymous):

hope you are unconfused now